# Advancing Security with Digital Twins: A Comprehensive Survey

### v1

@@ Abstract
The proliferation of electronic devices has greatly transformed every aspect of human life, such as communication, healthcare, transportation, and energy. Unfortunately, the global electronics supply chain is vulnerable to various attacks, including piracy of intellectual properties, tampering, counterfeiting, information leakage, side-channel, and fault injection attacks, due to the complex nature of electronic products and vulnerabilities present in them. Although numerous solutions have been proposed to address these threats, significant gaps remain, particularly in providing scalable and comprehensive protection against emerging attacks. Digital twin, a dynamic virtual replica of a physical system, has emerged as a promising solution to address these issues by providing backward traceability, end-to-end visibility, and continuous verification of component integrity and behavior. In this paper, we present a comprehensive survey of the application of digital twins based on their functional role and application domains. We comprehensively present the latest digital twin–based security implementations, including their role in cyber-physical systems, Internet of Things, cryptographic systems, detection of counterfeit electronics, intrusion detection, fault injection, and side-channel leakage. To the best of our knowledge, this is the first work to consider these critical security use cases within a single study to offer researchers and practitioners a unified reference for securing hardware with digital twins. Additionally, we discuss the integration of large language models into digital twins to enhance the security of systems by leveraging their advanced reasoning capabilities and highlight the challenges and limitations of applying digital twins to solve hardware security problems, with possible solutions. Further, we discuss current research trends and future directions in advancing hardware security through digital twins.
@@
Digital Twin, Heterogeneous Integration, IoT, Hardware Security, Large Language Models, Counterfeit ICs.
@@ I Introduction
The modern electronics supply chain is a highly intricate and globalized ecosystem that spans across multiple countries and involves various independent entities responsible for different stages of the product lifecycle, including design, fabrication, packaging, assembly, and distribution [1, 2]. This complexity is compounded by the presence of both trusted and untrusted actors, increasing the risk of vulnerabilities being introduced at any point in the supply chain [3]. The cross-border nature of operations makes oversight and accountability particularly challenging, as different regions may follow varying standards of security, regulation, and quality assurance. Furthermore, the supply chain consists of multiple hierarchical layers, often including subcontractors and third-party vendors, which makes it difficult for original equipment manufacturers to maintain complete visibility into every component’s origin and traceability information. A significant limitation is the lack of real-world implementation of a tracking system that provides the ability to trace components back to their source and verify their integrity [4, 5, 6]. This lack of transparency and visibility not only complicates risk management but also creates opportunities for adversaries to insert counterfeit parts, malicious circuitry, or compromise intellectual properties. As the demand for faster, cheaper, and more powerful electronics grows, addressing these challenges becomes critical to ensuring the security, reliability, and trustworthiness of electronic systems deployed across commercial, industrial, and national security applications.
@@
Although solutions exist to address these security challenges – such as counterfeit ICs [3, 2, 7], information leakage [8, 9, 10], hardware Trojans [11, 12, 13], and side-channel attacks [14, 15, 16]. However, there still exist research gaps as these approaches often focus on a specific threat, while the attack surface continues to expand due to the constant development of new threats by malicious actors [17, 18]. Therefore, there is a need for comprehensive and scalable solutions that can adapt to emerging threats. A promising solution to address these security concerns can be the adoption of digital twin (DT) technology, which creates a virtual replica of the physical object and enables real-time monitoring, forecasting, simulation, and analysis of each step in the manufacturing and delivery process.
@@
Leveraging these abilities, DTs are able to capture and analyze data throughout the lifecycle of electronics to detect anomalies indicative of counterfeit components. They support robust security analysis and real-time threat detection by continuously evaluating system behavior against expected virtual models. Additionally, they facilitate the simulation and emulation of physical systems within a virtual environment, allowing for the identification and mitigation of potential leakage points in designs early in the development phase to improve the security of the system. This paper, therefore, demonstrates the current status of digital twin-enabled hardware security by reviewing relevant research from recent years. There already exists a few work that surveys the use of digital twins for security [19, 20], but none is focused towards ensuring security in hardware (i.e., electronics). The main contribution of our work is to provide the history and fundamental concept of digital twins, a comprehensive review of DT applications in hardware security, and future research directions. Our paper makes the following contributions:
@@
  • Application of Digital Twins: We present detailed applications of digital twin technology across multiple domains, including manufacturing, smart cities, healthcare, and energy, highlighting how DTs enable real-time monitoring, predictive maintenance, performance optimization, and enhanced decision-making. Although few surveys exist in the literature [21, 19, 22], this survey focuses primarily on the security aspects.
@@
  • Digital Twin for Security Enhancement: We present a comprehensive survey of the latest digital twin–based security implementations and applications spanning both counterfeit detection and information‑leakage prevention in electronics. To the best of our knowledge, this is the first work to consider these critical security use cases within a single study to offer researchers and practitioners a unified reference for safeguarding hardware with digital twins. We also discuss the use of digital twins to enhance security in IoT and cyber-physical systems.
@@
  • Large Language Models in Digital Twins: The paper examines how the integration of large language models (LLMs) into digital twin frameworks strengthens hardware security via natural-language understanding and advanced reasoning capabilities. Such LLM-enhanced digital twins have the potential to enable conversational and interactive interfaces, generate detailed scene descriptions, and provide evidence-based decision-making. These features ensure secure and resilient operations across diverse applications, including hardware security.
@@
  • Future Directions: We outlined the current challenges and limitations associated with applying digital twins in hardware security, as identified in prior works [18, 23, 24, 25], along with proposed solutions to address them. Recent studies employing LLMs for security verification [26, 27] have demonstrated significant potential. Since digital twins and LLMs individually enhance security, we envision that their convergence will play a pivotal role in shaping the future of security solutions. Although the integration of DTs with LLMs is currently in the research phase within non-security domains, it is yet to be explored for security.
@@ II-B Concept of Digital Twins
The basic idea of the digital twin is to represent a physical product, system, or process in the virtual domain to monitor, simulate, and optimize its performance through real-time data integration. By doing so, we can obtain the same information from the virtual twin as could have been obtained if we had access to the physical object. In addition, we can perform what-if analysis on the virtual twin without putting the physical counterpart at risk. In principle, a digital twin consists of three (3) components, namely, the physical twin in the real space, the virtual twin, and the communication links between them [43, 44].
@@
Figure 1 shows a conceptual diagram of a digital twin [45]. The physical twin is the real-world product, process, system, or asset being digitally represented. It serves as the foundation or source for the digital twin’s data and functionality. This component of the digital twin resides in real space and interacts with its environment, performs its designated tasks, and undergoes wear and tear. All of these provide the essential data needed for the digital twin to simulate, analyze, and predict future behaviors and outcomes. On the other hand, there is the virtual or digital twin, which is a representation of the physical counterpart in the virtual space or environment, and is capable of emulating the behavior of its physical twin in great detail. The virtual twin can be modeled as a physics-based model from first principles. However, obtaining a physics-based model may be infeasible for some complex systems. In such situations where the physics-based models are not easily available, data-driven models are used. This approach is based on the assumption that data is a manifestation of both known and unknown physics [21]. To build this type of model, one is required to perform data collection, data pre-processing, and data analysis with Artificial Intelligence (AI). An example of this type of model is found in [37, 36]. Sometimes, a combination of physics-based and data-driven models is required to model the system. This type of digital twin model is referred to as the hybrid model [46, 47]. A graphical approach to modeling a system was discussed in [48]. In short, the virtual twin is the component of the digital twin that enables detailed simulations, optimization, diagnosis, and predictions to support decision-making. Lastly, the connection link is the component of digital twins that allows data to flow from the physical twin to the virtual twin in the form of sensor data and historical data. Depending on the system being modeled, information and decisions will flow from the virtual space to the real space.
@@
Digital twins are classified into three types [43] – digital twin prototype, digital twin instance, and digital twin aggregate, shown in the second ring of Figure 2 and described in Section II-C. The third ring describes the virtual twin modeling methods, which are grouped into physics-based, data-driven, hybrid, and graphical methods. In the fourth and fifth rings, we present the applications of digital twins based on their role and domain applied, respectively. These applications are described in Section III.
Figure 2: Digital Twin Taxonomy and Applications
@@ II-C Types of Digital Twins
  1. 1.
    Digital twin prototype (DTP): DTP represents a product in its pre-manufacturing stage. It includes all the necessary information to design, produce, and describe a physical counterpart, such as requirements, fully annotated 3D models, bill of materials (BoM), and bills of processes, services, and disposal.
  2. 2.
    Digital twin instance (DTI): DTI represents the digital counterpart of an individual manufactured product. It is created when the physical product rolls off the production line and is linked to its corresponding physical twin throughout its lifecycle. This connection allows for real-time synchronization and updates, enabling effective monitoring and management of the product’s performance, maintenance, and overall lifecycle.
  3. 3.
    Digital twin aggregate (DTA): DTA aggregates multiple DTIs, providing a composite view of all instances of a product or system. This aggregation facilitates large-scale analysis, such as identifying trends, optimizing fleet performance, and improving operational efficiency across multiple units.
@@ III Application of Digital Twin
In this section, we categorize the applications of digital twins and briefly discuss them along two dimensions: first, based on the functional role of the digital twin, and second, based on the specific domain in which it is applied.
@@ III-A Role-Based Application of DT
• Real-time Monitoring: This is the simplest function of a digital twin. Here, the digital twin simply reflects the state of the physical system at any point in time based on real-time data collected from the physical system. With this, no matter where the physical system is in the world, we can get information about its state and performance. For instance, we can monitor the health of industrial machines, bridges, vehicles, electronics, etc. A practical example is found in [49], where digital twin was used to monitor the Structural Health of bridges in Vietnam. The design incorporated sensors to collect operational data, and a fog layer for local pre-processing before transmitting the data to the cloud for further analysis and visualization. Similarly, Li et al. introduced a cloud-based Battery Management System (BMS) to build a digital twin of battery systems, ensuring accurate status monitoring [50]. Another application of digital twin for real-time monitoring can be found in [51], where Li et al. proposed a technique to monitor the health and performance of a boost converter in power electronics. In summary, as DT continues to evolve, its real-time monitoring applications will play an increasingly important role in optimizing performance and decision-making across various domains.
• Optimization: With digital twin, we have access to real-world data from the physical system, the virtual twin, and AI. The integration of these components for real-time simulations allows engineers to explore multiple design configurations and optimize critical parameters to improve the performance of the system. Let us consider a scenario such as the design of an aircraft. With the digital twin of the aircraft, we can simulate airflow, thermal stresses, and component interactions to refine the designs, leading to an optimal design that reduces drag and improves engine efficiency. Similarly, in manufacturing, digital twins have the potential to simulate material flow, machine-to-machine interactions, and energy consumption to optimize production layouts and waste. The application of digital twin for optimization can be found in  [52, 28], where the Digital Twin Shop Floor (DTS) was proposed to optimizes manufacturing through an iterative process that integrates real-time data from the Physical Shop Floor (PS), Virtual Shop Floor (VS), and Shop Floor Service System (SSS). The digital twin detects potential resource allocation conflicts and provides optimization strategies to enhance efficiency before actual execution. During manufacturing, PS transmits the real-time state data to VS, which updates itself to reflect the physical changes, ensuring alignment with predefined plans. This continuous feedback loop minimizes disruptions, enhances precision, and maximizes overall operational efficiency, making DTS a powerful tool for intelligent and adaptive manufacturing optimization.
@@
• Forecasting: Digital twin serves as a predictive tool to forecast the future behavior and performance of the physical twin. The digital twin achieves this function by using its AI component to analyze the observed measurements coming from sensors attached to the physical twin. With this, we could predict when a system or its subsystem will fail. Thereby, eliminating the unexpected failure of the system. It is important to know that this function of digital twins is the reason for the widespread adoption of digital twins in manufacturing in recent years, to reduce operational downtime. For example, in a recent study, Aivaliotis et al. demonstrated the use of digital twins to predict the health condition of a six-axis robotic structure used for welding in order to reduce downtime [54]. Similarly, the authors used digital twins to predict the health status and maintenance cycles of machines [55]. Similarly, [51, 50, 56] employed digital twins to predict the remaining useful life of systems.
• Diagnosis: Digital twin technology is a transformative leap in predictive maintenance and fault detection, enabling engineers to detect faults [57] and also trace the origin of such faults from the observed historical data of a physical system [18]. By employing advanced analytics, machine learning (ML), and real-time data integration, the diagnosis function of digital twins empowers industries to transition from reactive to proactive maintenance strategies. Ultimately, this improves systems’ reliability, sustainability, and performance.

### v2

@@ Abstract
The proliferation of electronic devices has greatly transformed every aspect of human life, such as communication, healthcare, transportation, and energy. Unfortunately, the global electronics supply chain is vulnerable to various attacks, including piracy of intellectual properties, tampering, counterfeiting, information leakage, side-channel, and fault injection attacks, due to the complex nature of electronic products and vulnerabilities present in them. Although numerous solutions have been proposed to address these threats, significant gaps remain, particularly in providing scalable and comprehensive protection against emerging attacks. Digital twin, a dynamic virtual replica of a physical system, has emerged as a promising solution to address these issues by providing backward traceability, end-to-end visibility, and continuous verification of component integrity and behavior. In this paper, we present a comprehensive survey of the application of digital twins based on their functional role and application domains. We comprehensively present the latest digital twin–based security implementations, including their role in cyber-physical systems, Internet of Things, cryptographic systems, detection of counterfeit electronics, intrusion detection, fault injection, and side-channel leakage. This work comprehensively considers these critical security use cases within a single study to offer researchers and practitioners a unified reference for securing hardware with digital twins. Additionally, we discuss the integration of large language models into digital twins to enhance the security of systems by leveraging their advanced reasoning capabilities and highlight the challenges and limitations of applying digital twins to solve hardware security problems, with possible solutions. Further, we discuss current research trends and future directions in advancing hardware security through digital twins.
@@
Digital Twin, IoT, Hardware Security, Large Language Models, Counterfeit ICs.
@@ I Introduction
The modern electronics supply chain is a highly intricate and globalized ecosystem that spans across multiple countries and involves various independent entities responsible for different stages of the product lifecycle, including design, fabrication, packaging, assembly, and distribution [1, 2]. This complexity is compounded by the presence of both trusted and untrusted actors, increasing the risk of vulnerabilities being introduced at any point in the supply chain [3]. The cross-border nature of operations makes oversight and accountability particularly challenging, as different regions may follow varying standards of security, regulation, and quality assurance. Furthermore, the supply chain consists of multiple hierarchical layers, often including subcontractors and third-party vendors, which makes it difficult for original equipment manufacturers to maintain complete visibility into every component’s origin and traceability information. A significant limitation is the lack of real-world implementation of a tracking system that provides the ability to trace components back to their source and verify their integrity [4, 5, 6]. This lack of transparency and visibility not only complicates risk management but also creates opportunities for adversaries to insert counterfeit parts, malicious circuitry, or compromise intellectual properties. As the demand for faster, cheaper, and more powerful electronics grows, addressing these challenges becomes critical to ensure the security, reliability, and trustworthiness of electronic systems deployed across commercial, industrial, and national security applications.
@@
Although solutions exist to address these security challenges – such as counterfeit ICs [3, 2, 7], information leakage [8, 9, 10], hardware Trojans [11, 12, 13], and side-channel attacks [14, 15, 16]. However, research gaps still exist as these approaches often focus on a specific threat, while the attack surface continues to expand due to the constant development of new threats by adversaries [17, 18]. Therefore, there is a need for comprehensive and scalable solutions for adapting to emerging threats. A promising solution to address these security concerns can be the adoption of digital twin (DT) technology, which creates a virtual replica of the physical object, providing real-time monitoring, forecasting, simulation, and analysis of each step in the manufacturing and delivery processes.
@@
With these abilities, DTs are able to capture and analyze data throughout the lifecycle of electronics to detect anomalies indicative of counterfeit components. They support robust security analysis and real-time threat detection by continuously evaluating system behavior against expected virtual models. In addition, they facilitate the simulation and emulation of physical systems within a virtual environment, allowing for the identification and mitigation of potential leakage points in designs early in the development phase to improve the security of the system. This paper, therefore, demonstrates the current status of digital twin-enabled hardware security by reviewing relevant research from recent years. To contextualize this survey within existing literature, a comparative overview of related works is provided in Table I. While prior surveys have focused on specific domains such as CPS [19, 20], or IoT applications [21]. In contrast, this survey offers a broader and unified perspective by covering CPS, IoT, supply chain, and cryptographic systems in a single study, while also examining the integration of LLMs with DT for intelligent security analysis.
Table I: Positioning this survey among related works.
Year Ref. Domains Covered
IoT CPS Supply Chain Crypto Systems LLM
2019  [19] $\checkmark$
2020  [20] $\checkmark$
2024  [21] $\checkmark$ $\checkmark$
This Study $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
The main contribution of our work is to provide the history and fundamental concept of digital twins, a comprehensive review of DT applications in hardware security, and future research directions. Key contributions of this work include:
@@
•  Application of Digital Twins: We present detailed applications of DT across multiple domains, including manufacturing, smart cities, healthcare, and energy, highlighting how DTs enable real-time monitoring, predictive maintenance, performance optimization, and enhanced decision-making. Although few surveys exist in the literature [22, 20, 23], this survey focuses primarily on the security aspects.
@@
•  Digital Twin for Security Enhancement: We present a comprehensive survey of the latest DT–based security implementations and applications spanning both counterfeit detection and information‑leakage prevention in electronics. This work comprehensively considers critical security use cases within a single study to offer researchers and practitioners a unified reference for safeguarding electronic hardware with digital twins. We also discuss the use of DTs to enhance security in IoT and cyber-physical systems.
@@
•  Large Language Models in Digital Twins: The paper examines how the integration of large language models (LLMs) into DT frameworks strengthens hardware security via natural-language understanding and advanced reasoning. Such LLM-enhanced DTs have the potential to enable conversational and interactive interfaces, generate detailed scene descriptions, and provide evidence-based decision-making. These features ensure secure and resilient operations across diverse applications, including hardware security.
@@
•  Future Directions: We outlined the current challenges and limitations associated with applying DTs in hardware security, as identified in prior works [18, 24, 25, 26], along with proposed solutions to address them. Recent studies employing LLMs for security verification [27, 28] have demonstrated significant potential. Since DTs and LLMs individually enhance security, we envision that their convergence will play a pivotal role in shaping the future of security solutions. Although the integration of DTs with LLMs is currently in the research phase within non-security domains, it has yet to be explored for security.
@@ II-B Concept of Digital Twins
The basic idea of the digital twin is to represent a physical product, system, or process in the virtual domain to monitor, simulate, and optimize its performance through real-time data integration. By doing so, we can obtain the same information from the virtual twin as could have been obtained if we had access to the physical object. In addition, we can perform what-if analysis on the virtual twin without putting the physical counterpart at risk. In principle, a digital twin consists of three (3) components, namely, the physical twin in the real space, the virtual twin, and the communication links between them [44, 45]. Figure 1 shows a conceptual diagram of a digital twin [46].
@@
The physical twin is the real-world product, process, system, or asset being digitally represented. It serves as the foundation or source for the digital twin’s data and functionality. This component of the digital twin resides in real space and interacts with its environment, performs its designated tasks, and undergoes wear and tear. All of these provide the essential data needed for the digital twin to simulate, analyze, and predict future behaviors and outcomes. On the other hand, there is the virtual or digital twin, which is a representation of the physical counterpart in the virtual space or environment, and is capable of emulating the behavior of its physical twin in great detail. The virtual twin can be modeled as a physics-based model from first principles. However, obtaining a physics-based model may be infeasible for some complex systems. In such situations where the physics-based models are not easily available, data-driven models are used. This approach is based on the assumption that data is a manifestation of both known and unknown physics [22]. To build this type of model, one is required to perform data collection, data pre-processing, and data analysis with Artificial Intelligence (AI). An example of this type of model is found in [38, 37]. Sometimes, a combination of physics-based and data-driven models is required to model the system. This type of digital twin model is referred to as the hybrid model [47, 48]. A graphical approach to modeling a system was discussed in [49]. In short, the virtual twin is the component of the digital twin that enables detailed simulations, optimization, diagnosis, and predictions to support decision-making. Lastly, the connection link is the component of digital twins that allows data to flow from the physical twin to the virtual twin in the form of sensor data and historical data. Depending on the system being modeled, information and decisions will flow from the virtual space to the real space.
Figure 2: Digital Twin Taxonomy and Applications.
@@
Digital twins are classified into three types [44] – digital twin prototype, digital twin instance, and digital twin aggregate, shown in the second ring of Figure 2 and described in Section II-C. The third ring describes the virtual twin modeling methods, which are grouped into physics-based, data-driven, hybrid, and graphical methods. In the fourth and fifth rings, we present the applications of digital twins based on their role and the domain applied, respectively. These applications are described in Section III.
@@ II-C Types of Digital Twins
•  Digital twin prototype (DTP) represents a product in its pre-manufacturing stage. It includes all the necessary information to design, produce, and describe a physical counterpart, such as requirements, fully annotated 3D models, bill of materials (BoM), and bills of processes, services, and disposal.
•  Digital twin instance (DTI) represents the digital counterpart of an individual manufactured product. It is created when the physical product rolls off the production line and is linked to its corresponding physical twin throughout its lifecycle. This connection allows for real-time synchronization and updates, enabling effective monitoring and management of the product’s performance, maintenance, and overall lifecycle.
•  Digital twin aggregate (DTA) aggregates multiple DTIs, providing a composite view of all instances of a product or system. This aggregation facilitates large-scale analysis, such as identifying trends, optimizing fleet performance, and improving operational efficiency across multiple units.
@@ III Application of Digital Twin
In this section, we categorize the applications of digital twins and briefly discuss them along two dimensions: first, based on the functional role of the digital twin, and second, based on the domain in which it is applied.
@@ III-A Role-Based Application of DT
• Real-time Monitoring: This is the simplest function of a digital twin. Here, the digital twin simply reflects the state of the physical system at any point in time based on real-time data collected from the physical system. With this, no matter where the physical system is in the world, we can get information about its state and performance. For instance, we can monitor the health of industrial machines, bridges, vehicles, electronics, etc. A practical example is found in [50], where a digital twin was used to monitor the Structural Health of bridges in Vietnam. The design incorporated sensors to collect operational data, and a fog layer for local pre-processing before transmitting the data to the cloud for further analysis and visualization. Similarly, Li et al. introduced a cloud-based Battery Management System (BMS) to build a digital twin of battery systems, ensuring accurate status monitoring [51]. Another application of digital twin for real-time monitoring can be found in [52], where Li et al. proposed a technique to monitor the health and performance of a boost converter in power electronics. In summary, as DT continues to evolve, its real-time monitoring applications will play an increasingly important role in optimizing performance and decision-making across various domains.
• Optimization: With digital twin, we have access to real-world data from the physical system, the virtual twin, and AI. The integration of these components for real-time simulations allows engineers to explore multiple design configurations and optimize critical parameters to improve the performance of the system. Let us consider a scenario such as the design of an aircraft. With the digital twin of the aircraft, we can simulate airflow, thermal stresses, and component interactions to refine the designs, leading to an optimal design that reduces drag and improves engine efficiency. Similarly, in manufacturing, digital twins have the potential to simulate material flow, machine-to-machine interactions, and energy consumption to optimize production layouts and waste. The application of digital twin for optimization can be found in  [53, 29], where the Digital Twin Shop Floor (DTS) was proposed to optimize manufacturing through an iterative process that integrates real-time data from the Physical Shop Floor (PS), Virtual Shop Floor (VS), and Shop Floor Service System (SSS). The digital twin detects potential resource allocation conflicts and provides optimization strategies to enhance efficiency before actual execution. During manufacturing, PS transmits the real-time state data to VS, which updates itself to reflect the physical changes, ensuring alignment with predefined plans. This continuous feedback loop minimizes disruptions, enhances precision, and maximizes overall operational efficiency, making DTS a powerful tool for intelligent and adaptive manufacturing optimization.
@@
• Forecasting: Digital twin serves as a predictive tool to forecast the future behavior and performance of the physical twin. The digital twin achieves this function by using its AI component to analyze the observed measurements coming from sensors attached to the physical twin. With this, we could predict when a system or its subsystem will fail. Thereby eliminates the unexpected failure of the system. It is important to know that this function of digital twins is the reason for the widespread adoption of digital twins in manufacturing in recent years, to reduce operational downtime. For example, in a recent study, Aivaliotis et al. demonstrated the use of digital twins to predict the health condition of a six-axis robotic structure used for welding in order to reduce downtime [55]. Similarly, the authors used digital twins to predict the health status and maintenance cycles of machines [56]. Similarly, [52, 51, 57] employed digital twins to predict the remaining useful life of systems.
• Diagnosis: Digital twin technology is a transformative leap in predictive maintenance and fault detection, enabling engineers to detect faults [58] and also trace the origin of such faults from the observed historical data of a physical system [18]. By employing advanced analytics, machine learning (ML), and real-time data integration, the diagnosis function of digital twins empowers industries to transition from reactive to proactive maintenance strategies. Ultimately, improves systems’ reliability, sustainability, and performance.